OpenStudy (exoexoexo):

toko follows a budget to buy supplies to paint his room. he spends 1/2 of his budget on a roller, a tray , and tape. he spends $12.27 on paint, and of the remaining money, he spends 2/3 on a trim brush, which costs $3.82. how much money has he budgeted

pooja195 (pooja195):

\[ \huge~\rm~b / 2 = $12.27 + [$3.82 / (2/3)]\] Solve this

pooja195 (pooja195):

We need to simplfy the expression it simplifies to \[\huge~\rm~\frac{ 1 }{ 2 }b=18\]

pooja195 (pooja195):

btw b=budget

pooja195 (pooja195):

From here you multiply be the RECIPROCAL(flipped fraction) \[\huge~\rm~(2)\frac{ 1 }{ 2 }b=18(2)\]

pooja195 (pooja195):

Im sure you can solve from here c:

OpenStudy (exoexoexo):

b=36

pooja195 (pooja195):

good :)

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)